The average height of all entering freshmen students is 165 cm with the standard deviation


. To test the claim, a researcher selected a random sample of 50 freshmen of a certain


college. In this sample, and . Is the claim true?

1
Expert's answer
2022-05-31T12:13:42-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=165"

"H_1:\mu\not=165"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c =1.96."

The rejection region for this two-tailed test is "R = \{z:|z|> 1.96\}."

The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{163-165}{7/\sqrt{50}}=-2.0203"


6. Since it is observed that "|z|=2.0203>1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed is "p=2P(Z<-2.0203)=0.043352," and since "p=0.043352<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 300, at the "\alpha = 0.05" significance level.